OpenStudy (anonymous):

substitution method 8x-4y=36 y=2x+9

OpenStudy (anonymous):

This question must be incorrect, can you check if you copied it down correctly?

OpenStudy (mimi_x3):

8x - 4(2x+9) = 36 8x - 8x - 36 = 36 lols , the x vanishes

OpenStudy (mimi_x3):

so no solution

OpenStudy (angela210793):

replace y in the 1st eq 8x-4(2x+9)-36=0 8x-8x-36-36=0 -72=0????? not true....as Camp said check if u've written it right or may be no solutions

OpenStudy (anonymous):

thanks. I did write it down correctly so there is no solution

OpenStudy (angela210793):

ok :)